40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Oct 8 08:08:18 2019, gautam, Update, PEM, PEM BLRMS anomaly PEManomaly.png
    Reply  Wed Oct 23 11:49:21 2019, gautam, Update, PEM, PEM BLRMS anomaly seisAll_20191021.pdfspecGrams.zip
       Reply  Thu Oct 24 18:37:15 2019, gautam, Update, PEM, T240 checkout 
          Reply  Tue Nov 5 12:37:50 2019, gautam, Update, PEM, T240 interface unit pulled out T240_recenter.pdfffPotential.pdf
             Reply  Wed Nov 13 19:34:45 2019, gautam, Update, PEM, Follow-up on seismometer discussion seisAll_20191111.pdfffPotential.pdf
                Reply  Wed Nov 13 20:15:56 2019, rana, Update, PEM, Follow-up on seismometer discussion 
                   Reply  Wed Nov 13 23:40:15 2019, gautam, Update, PEM, Follow-up on seismometer discussion seisAll_20191111_1Hz.pdf
                      Reply  Thu Nov 14 12:11:04 2019, rana, Update, PEM, Follow-up on seismometer discussion 
                      Reply  Fri Nov 15 00:18:41 2019, rana, Update, PEM, Follow-up on seismometer discussion MCshake.pdf
                         Reply  Fri Nov 15 12:16:48 2019, gautam, Update, PEM, Follow-up on seismometer discussion BS_ZspecGram.pdf.zip
                            Reply  Mon Nov 18 14:32:53 2019, gautam, Update, PEM, Follow-up on seismometer discussion 
                               Reply  Tue Nov 19 21:53:57 2019, gautam, Update, PEM, Follow-up on seismometer discussion seisxSpec.pdf
Message ID: 14949     Entry time: Tue Oct 8 08:08:18 2019     Reply to this: 14989
Author: gautam 
Type: Update 
Category: PEM 
Subject: PEM BLRMS anomaly 

Yesterday, Koji and I noticed (from the wall StripTool traces) that the vertex seismometer RMS between 0.1-0.3 Hz in the X-direction increased abruptly around 6pm PDT. This morning, when I came in, I noticed that the level had settled back to the normal level. Trending the BLRMS channels over the last 24 hours, I  see that the 0.3-1 Hz band in the Z direction shows some anomalous behaviour almost in the exact same time-band. Hard to believe that any physical noise was so well aligned to the seismometer axes, I'm inclined to think this is indicative of some electronics issues with the Trillium interface unit, which has been known to be flaky in the past.

Attachment 1: PEManomaly.png  54 kB  Uploaded Tue Oct 8 09:14:06 2019  | Hide | Hide all
PEManomaly.png
ELOG V3.1.3-